OpenStudy (anonymous):

solve and show all work plz y = x2 + 3x + 2

OpenStudy (anonymous):

Solve for what?

OpenStudy (anonymous):

its a quadratic equation so :/

OpenStudy (anonymous):

y = x^2 + 3x + 2 Factor the equation. y = (x+2)(x+1) Set the equation equal to zero. y = (x+2)(x+1) = 0 x+2 = 0 x = -2 x+1 = 0 x = -1 Therefore, x = -1 and x = -2.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Oh, you mean solve for y = 0. You should be more specific. You could just as easily solve for y = \(\pi\) or x = 15, or anything really.